Grasping the Concept

Use the following thermochemical equations as needed to find the heat of formation of diamond:

C(diamond) + O2(g) → CO2(g) ΔHo = -395.3 KJ

2 CO2(g) → 2 CO(g) + O2(g) ΔHo= 566.0 KJ

C(graphite) + O2(g) → CO2(g) ΔHo = -395.3 KJ

2 CO(g) → C(graphite) + CO2(g) ΔHo= -172.5 KJ


1
Expert's answer
2021-10-26T05:38:35-0400

eq - equation

(1st eq): C(diamond) + O2(g) → CO2(g), ΔHo = −395.3 kJ

(2nd eq): 2CO2(g) → 2CO(g) + O2(g), ΔHo = 566.0 kJ

(3d eq): C(graphite) + O2(g) → CO2(g), ΔHo = −393.5 kJ (NB! −393.5 kJ not −395.3 kJ)

(4th eq): 2CO(g) → C(graphite) + CO2(g), ΔHo = −172.5 kJ


C(graphite) → C(diamond), ΔHx = ???


Solution:

According to Hess's law, the heat of reaction depends upon initial and final conditions of reactants and does not depend of the intermediate path of the reaction.


1) Modify the four given equations to get the target equation:

(1st eq): multiply by −1

(2nd eq): do nothing

(3d eq): multiply by 2

(4th eq): do nothing


2) Rewrite four equations with the changes made (including changes in their enthalpy):

(1*st eq): CO2(g) → C(diamond) + O2(g), ΔH1 = (−1) × (−395.3 kJ) = 395.3 kJ

(2nd eq): 2CO2(g) → 2CO(g) + O2(g), ΔH2 = 566.0 kJ

(3*d eq): 2C(graphite) + 2O2(g) → 2CO2(g), ΔH3 = (2) × (−393.5 kJ) = −787.0 kJ

(4th eq): 2CO(g) → C(graphite) + CO2(g), ΔH4 = −172.5 kJ


3) Cancel out the common species on both sides:

CO2(g) ⇒ the sum of 1*st and 2nd equations versus the sum of 3*d and 4th equations

O2(g) ⇒ 3*d equation versus the sum of 1st and 2nd equations

CO(g) ⇒ 4th equation versus 2nd equation


Thus, adding modified equations and canceling out the common species on both sides, we get:

C(graphite) → C(diamond)


4) Add the ΔH values of (1*st), (2nd), (3*d) and (4th) equations to get your answer:

ΔHx = ΔH1 + ΔH2 + ΔH3 + ΔH4

ΔHx = (395.3 kJ) + (566.0 kJ) + (−787.0 kJ) + (−172.5 kJ) = 1.8 kJ

ΔHx = 1.8 kJ


C(graphite) → C(diamond), ΔHo = 1.8 kJ


Answer: ΔHo = 1.8 kJ